4. Complete the following: a. 26 Fe +→žn+ 51 b. __+y +n + Be
Answer #1

- Liet, the missing nuclide is my Now, 26+2 = 0+A = A=28 :9 54+4 = 1+M + M=52-1=57 teste pet god — nu se m + 8 intable 1. Le